What is the repulsive electrical force between two protons in a nucleus that are 4.6x 10-l5 m apart from each other?
n.b. give your answer in Newtons
Answer #1

Electric force is given by the Coulomb's law

F = k* q1*q2/r​​​​​​2

where ,

K = 9* 109

q1 and q1 are charges in coulomb

r = distance between charges

charge of proton = 1.6* 10-19

r = 4.6*10-15

Therefore ,

F = 9*109 *1.6* 10-19 *1.6*10-19/(4.6*10-15)2

F= 1.08*101 N = 10.8 N